Bulluck Corporation makes a product with the following standard costs:

Standard Quantity or Hours Standard Price or Rate
Direct materials 3.5 grams $ 1.00 per gram
Direct labor 0.7 hours $ 11.00 per hour
Variable overhead 0.7 hours $ 2.00 per hour

The company reported the following results concerning this product in July.

Actual output 3,000 units
Raw materials used in production 11,370 grams
Actual direct labor-hours 1,910 hours
Purchases of raw materials 12,100 grams
Actual price of raw materials purchased $ 1.20 per gram
Actual direct labor rate $ 11.40 per hour
Actual variable overhead rate $ 2.10 per hour

The company applies variable overhead on the basis of direct labor-hours. The direct materials purchases variance is computed when the materials are purchased.

The variable overhead rate variance for July is:

Multiple Choice

  • $191 U

  • $210 U

  • $191 F

  • $210 F

The Tipton Division of Dudley Company reported the following data last year:

Return on investment 20 %
Minimum required rate of return 12 %
Residual income $ 50,000

Tipton Division's average operating assets last year were:

Multiple Choice

  • $250,000

  • $625,000

  • $333,333

  • $416,677

Kuzio Corporation produces and sells a single product. Data concerning that product appear below:

Per Unit Percent of Sales
Selling price $ 150 100 %
Variable expenses 75 50 %
Contribution margin $ 75 50 %

The company is currently selling 6,500 units per month. Fixed expenses are $206,000 per month. The marketing manager believes that a $6,300 increase in the monthly advertising budget would result in a 100 unit increase in monthly sales. What should be the overall effect on the company's monthly net operating income of this change?

rev: 03_09_2018_QC_CS-121313

Multiple Choice

  • increase of $7,500

  • increase of $1,200

  • decrease of $6,300

  • decrease of $1,200

The Assembly Department started the month with 24,000 units in its beginning work in process inventory. An additional 309,000 units were transferred in from the prior department during the month to begin processing in the Assembly Department. There were 29,000 units in the ending work in process inventory of the Assembly Department. How many units were transferred to the next processing department during the month?

Multiple Choice

  • 362,000

  • 333,000

  • 304,000

  • 314,000

Bonilla Inc. has a $700,000 investment opportunity with the following characteristics:

Sales $ 2,240,000
Contribution margin ratio 40 % of sales
Fixed expenses $ 739,200

The ROI for the investment opportunity is closest to:

Multiple Choice

  • 22.4%

  • 21.2%

  • 128.0%

  • 7.0%

Data concerning three of the activity cost pools of Salcido LLC, a legal firm, have been provided below:

Activity Cost Pool Total Cost Total Activity
Researching legal issues $ 21,080 680 research hours
Meeting with clients $ 1,195,279 7,333 meeting hours
Preparing documents $ 92,440 5,780 documents

The activity rate for the “meeting with clients” activity cost pool is closest to:

Multiple Choice

  • $72 per meeting hour

  • $91 per meeting hour

  • $1,195,279 per meeting hour

  • $163 per meeting hour
Answer #1

1) Variable overhead rate variance = (standard rate-actual rate)actual hours = (2-2.10)*1910 = 191 U

So answer is a) $191 U

2) Average operating assets = 50000*100/8 = 625000

So answer is b) $625000

3) Incremental profit (loss) = Incremental revenue-Incremental cost = (100*75-6300) = 1200

So answer is b) Increase of $1200

4) Unit transferred out = Beginning WIP+Unit started-ending WIP = 24000+309000-29000 = 304000

So answer is c) 304000

5) ROI = Operating income/Average operating assets = 156800/700000 = 22.4%

So answer is a) 22.4%

6) Activity rate = 1195279/7333 = 163 per meeting hour

So answer is d) $163 per meeting hour